# Face Forgery Detection with Elaborate Backbone

### v1

@@ Abstract
Face Forgery Detection (FFD), or Deepfake detection, aims to determine whether a digital face is real or fake. Due to different face synthesis algorithms with diverse forgery patterns, FFD models often overfit specific patterns in training datasets, resulting in poor generalization to other unseen forgeries. This severe challenge requires FFD models to possess strong capabilities in representing complex facial features and extracting subtle forgery cues. Although previous FFD models directly employ existing backbones to represent and extract facial forgery cues, the critical role of backbones is often overlooked, particularly as their knowledge and capabilities are insufficient to address FFD challenges, inevitably limiting generalization. Therefore, it is essential to integrate the backbone pre-training configurations and seek practical solutions by revisiting the complete FFD workflow, from backbone pre-training and fine-tuning to inference of discriminant results. Specifically, we analyze the crucial contributions of backbones with different configurations in FFD task and propose leveraging the ViT network with self-supervised learning on real-face datasets to pre-train a backbone, equipping it with superior facial representation capabilities. We then build a competitive backbone fine-tuning framework that strengthens the backbone’s ability to extract diverse forgery cues within a competitive learning mechanism. Moreover, we devise a threshold optimization mechanism that utilizes prediction confidence to improve the inference reliability. Comprehensive experiments demonstrate that our FFD model with the elaborate backbone achieves excellent performance in FFD and extra face-related tasks, i.e., presentation attack detection. Code and models are available at https://github.com/zhenglab/FFDBackbone.
@@ 1 Introduction
HUMAN faces, being unique and readily available biometric data, play a pivotal role in various research fields, including face recognition [1, 2], generation [3, 4], and analysis [5, 6]. Unfortunately, facial information leakage and theft have spawned a clandestine industry where criminals exploit forged faces to perpetrate illegal activities, posing an extreme threat to national security, social stability, and individual well-being. Meanwhile, diverse GAN-based and Diffusion Model-based methods [7, 8, 9] in Artificial Intelligence Generated Content (AIGC) have drastically reduced the barriers to creating lifelike forged faces. The proliferation of these forged faces online has led to significant crises in the security and authenticity of information. Therefore, distinguishing digitally forged faces created by image/video synthesis algorithms, i.e., Face Forgery Detection (FFD), also referred to as Deepfake detection, has become a critical issue in both research and industrial communities [10, 11, 12].
Due to the diverse forgery patterns induced by different types of synthesis algorithms (e.g., entire face synthesis [13], face swapping [14], face attribute editing [15], and face reenactment [16]), FFD models often overfit specific forgery patterns within training datasets while neglecting implicit forgery cues shared across datasets, severely limiting their generalization. Unlike general image classification tasks that mainly concentrate on salient objects, FFD is a weak signal detection and classification task, where the increasing realism of forged faces obscures forgery cues behind complex facial features, making it more challenging for FFD models to uncover implicit forgery cues. Consequently, FFD models with better generalization must excel in two critical capabilities: robust facial feature representation and precise weak-forgery-cue extraction. Notably, both critical capabilities essentially stem from pre-training and fine-tuning of the backbone as a foundational model, which is a rational paradigm for improving generalization and achieving remarkable performance in various CV and NLP tasks. However, previous FFD methods often overlook the backbone’s capabilities, leading to unsatisfactory generalization.
Indeed, previous FFD methods also employ existing backbones as a fundamental unit in their models and propose various fine-tuning techniques to equip the backbone for better generalization. However, these backbones are typically used directly without adequate attention, especially regarding the compatible relationship between their differing capabilities and FFD challenges. Furthermore, we are surprised that most backbones used in FFD models are CNN-based [17, 18, 19, 20] rather than Transformer-based [21] networks, all of which are pre-trained with supervised learning (SL) on ImageNet for general scene tasks such as natural image classification and recognition. In contrast, backbones with self-supervised learning (SSL) and Transformer-based networks have proven superior feature representation capabilities in numerous CV tasks [22, 23], with their importance also highlighted by foundation models like MAE [24], BEiT [25, 26], CLIP [27], and GPT [28]. We reason that different backbone configurations can significantly influence the sensitivity of FFD models to subtle anomalous signals.
[…]
@@
Figure 1: Quantitative comparison of FFD methods with various backbones. “Unpre-trained” means networks with randomly initialized parameters, while “SL-1K” denotes those pre-trained with supervised learning on ImageNet-1K. The average AUC across three cross-datasets (Celeb-DF, DFDC, and FFIW) better reflects the generalization of FFD models.
It is well known that developing a highly generalized backbone necessitates massive datasets as the knowledge source, high-capacity networks as the carrier, and promising training methods as the drivers. For the datasets, facial images exhibit significant deviations from generic natural images (e.g., ImageNet) due to their unique inherent characteristics, including specific shapes, key point distributions, and locally consistent displacement fields [34, 35]. However, most backbones used in FFD models are pre-trained on ImageNet, which insufficiently represents fine-grained facial features. Indeed, in order to recognize real faces and capture their subtle differences from forged faces, a sophisticated FFD backbone may need three levels of knowledge: the general knowledge of image representation from large-scale natural images, the specific knowledge of facial representation from real faces, and the expert knowledge of identifying forgery cues from forged faces. For the networks, as shown in Figure 2, forgery clues often appear in local regions and their global differences within the face, requiring the networks to carry extensive facial knowledge and better capture long-range correlations and differences. For the training methods, as the core bridges that assemble knowledge onto the carrier, their pre-training and fine-tuning methods must closely match facial characteristics and FFD challenges, especially in exploring implicit forgery cues within facial components (e.g., eyebrows, eyes, and lips). Building upon this, we strive to develop an elaborate FFD backbone by delving into the correspondent relationship between the backbone’s capabilities and FFD problems.
Given that the final performance of FFD methods heavily depends on backbones, equipping the elaborate backbone with general, specific, and expert knowledge in the pre-training and fine-tuning stages is a rational and promising solution to enhancing the generalization of FFD models. Moreover, FFD’s ultimate goal is to determine whether the input face is real or fake, but FFD models usually output continuous prediction probabilities ($p\in[0,1]$) obtained through the Softmax function. Existing works empirically use $0.5$ as the threshold, converting continuous probabilities into discrete results (real or fake). However, it is hard to say that $0.5$ is the optimal threshold, as it varies significantly across different datasets and FFD models, and Softmax often produces overconfident prediction scores, which can lead to severe consequences in cases of overconfident incorrect predictions [36, 37]. Therefore, we seek a threshold optimization mechanism by integrating prediction probability and its confidence during the inference stage, further improving the backbone’s reliability and practicality.
@@
In addition to digitally forged faces, a category of fake faces made with physical materials, such as printed photos, replayed images/videos, and 3D masks, which also pose a significant threat in real-life scenarios, particularly the attacks on face recognition systems. Research on detecting these attacks, known as presentation attack detection [38, 39, 40] (PAD), face anti-spoofing [41, 42, 43, 44, 45, 46, 47], and face liveness detection [48, 49, 50], has also garnered widespread attention [51]. Since FFD and PAD are both binary classification tasks for distinguishing real or fake faces, we consider that the backbones can still play an essential role in PAD. Therefore, we further apply our FFD method to PAD to validate its potential and broaden its applicability.
@@
In summary, the main contributions of this work are as follows: (1) we propose a revitalized FFD pipeline through a complete revisiting from backbone pre-training and fine-tuning to inference of discriminant results; (2) we develop and pre-train a proficient backbone that better represents facial component features tailored for FFD task; (3) we build a competitive backbone fine-tuning framework to enhance the backbone’s capabilities in extracting diverse forgery cues; (4) we devise a threshold optimization mechanism to improve the backbone’s accuracy and reliability in inference results; (5) we conduct extensive experiments to demonstrate the superiority and potential of our elaborate backbone, achieving significantly better generalization than previous methods in FFD and extra face-related PAD tasks.
@@ 2 Related Work
Our work strives to seek insightful guidance and practical solutions by comprehensively revisiting FFD workflow and analyzing the critical role of backbones with different pre-training configurations in FFD task. We start by summarizing existing FFD methods and then describe supervised and self-supervised learning, both of which enhance the feature representation capabilities of backbones.
@@ 2.1 Face Forgery Detection
Thanks to the creation of numerous face forgery datasets such as FaceForensic++ [52], Celeb-DF [53], and DFDC [54], deep learning-based FFD methods have made considerable advancements and can achieve outstanding accuracy within training datasets (exceeding $95$%). However, their performance drops sharply when encountering unseen forgery patterns across different datasets. Recent FFD work aims to improve model generalization by implementing specific fine-tuning methods such as data augmentation [55, 56, 57, 20], attention mechanisms [58, 59, 60, 61, 62, 63, 21], forgery feature decoupling [64, 65, 66, 67, 68, 69, 70, 17, 71], and temporal inconsistency [72, 73, 57, 19], as well as spatial and frequency fusion [74, 75, 76].
@@
Due to the presence of redundant forgery-independent features in forged faces (e.g., identity, natural artifacts), the distinction of forgery-dependent features is disrupted and hindered. Researchers leveraged disentangled representation learning [78] or game-theoretical approaches to enhance the backbone’s ability for extracting forgery clues. Specifically, disentangled representation learning-based FFD methods primarily adopted a training process that involved exchanging decoupled forgery-independent and forgery-dependent features to reconstruct pairs between real and forged faces, enhancing the discriminative performance of the classification branch with the backbone encoder [64, 65, 66, 67, 69, 68, 17]. Game-theoretical-based FFD methods [79, 80] utilized the knowledge of game theory to decouple unrelated regions from source and target images, promoting FFD performance with the support of Shapley value [81].
FTCN [72] and AltFreezing [57] explored temporal inconsistencies in forged videos. FTCN reduced the spatial convolution kernel, while AltFreezing alternated between spatial and temporal freezing, both aiming to encourage ResNet3D to explore generic features of forged face videos. Choi et al[19] employed StyleGRU and a style attention module to exploit the temporal low-variance properties of style latent vectors in each frame for detecting forged videos. Additionally, Wang et al[74] and Miao et al[75] leveraged frequency domain information to assist the backbone in capturing more fine-grained artifact features.
After reviewing previous FFD methods and their technical implementations, we find that these methods consistently equip various fine-tuning techniques to backbones, such as ResNet [29], Xception [30], EfficientNet [31], and ViT [32], which are pre-trained with supervised learning on ImageNet. These backbones are often used directly in the fine-tuning stage while overlooking the crucial roles of both the pre-training and inference stages, as shown in Figure 4(top). In contrast, we endeavor to offer comprehensive research insights and guidance for FFD generalization by revisiting the complete FFD workflow.
@@ 2.2 Supervised and Self-Supervised Learning
Early backbone research utilized supervised learning (SL) with an image classification task as the pre-training approach, focusing on designing more sophisticated network architectures to improve backbone capacity and capability. During this period, numerous well-designed networks developed for the ImageNet Large-Scale Visual Recognition Challenge [82] demonstrated outstanding performance, such as GoogLeNet [83], VGG [84], ResNet [29], and SENet [85]. Benefiting from Transformers’ global contextual attention design, backbone network architectures have gradually transitioned from CNNs to Transformers. In particular, Vision Transformers (ViTs) [32] have come to dominate the field of computer vision (CV), achieving superior performance compared to ResNet, Xception, and EfficientNet.
@@
Given the significance of learning approaches with different pretext tasks for backbones in the pre-training stage, we conduct a comprehensive empirical study to evaluate their performance in FFD task and provide valuable guidelines for their selection and optimization.
@@ 3 FFD Workflow and Method
Figure 3: Overview of our research to revisit the complete FFD workflow from backbone pre-training and fine-tuning to inference in discriminant results. This includes network architectures, datasets, learning approaches with supervised learning (SL) and self-supervised learning (SSL), which encompass contrastive learning (CL) and masked image modeling (MIM) pretext tasks, as well as our competitive backbone fine-tuning framework and threshold optimization mechanism.
@@ 3.1 Research Overview
With the continuous expansion of massive data and computational resources, deep learning-based works increasingly follow the three-stage pipeline of pre-training, fine-tuning, and inference. As a typical binary classification task, we consider that the build process of FFD models also adheres to this pipeline, although these existing works primarily focus on optimizing fine-tuning techniques. Therefore, we aim to revisit the FFD workflow from backbone development to final applications and further develop a practical and reliable backbone for FFD task, which ultimately determines whether an input face is real or fake.
Backbone serves as the starting point and foundation of FFD workflow, and following our investigation of backbones with different pre-training configurations, we construct a comprehensive research overview, as shown in Figure 3. This overview covers the complete FFD workflow: (1) the network architectures, (2) the learning approaches with SL and SSL, as well as the datasets such as ImageNet and real faces during pre-training, (3) the adaptation methods to downstream tasks during fine-tuning, and (4) the application mechanisms to maximize performance during inference.

### v2

# Revisiting Face Forgery Detection: From Facial Representation to Forgery Detection

@@ Abstract
Face Forgery Detection (FFD), or Deepfake detection, aims to determine whether a digital face is real or fake. Due to different face synthesis algorithms with diverse forgery patterns, FFD models often overfit specific patterns in training datasets, resulting in poor generalization to other unseen forgeries. Existing FFD methods primarily leverage pre-trained backbones with general image representation capabilities and fine-tune them to identify facial forgery cues. However, these backbones lack domain-specific facial knowledge and insufficiently capture complex facial features, thus hindering effective implicit forgery cue identification and limiting generalization. Therefore, it is essential to revisit FFD workflow across the pre-training and fine-tuning stages, achieving an elaborate integration from facial representation to forgery detection to improve generalization. Specifically, we develop an FFD-specific pre-trained backbone with superior facial representation capabilities through self-supervised pre-training on real faces. We then propose a competitive fine-tuning framework that stimulates the backbone to identify implicit forgery cues through a competitive learning mechanism. Moreover, we devise a threshold optimization mechanism that utilizes prediction confidence to improve the inference reliability. Comprehensive experiments demonstrate that our method achieves excellent performance in FFD and extra face-related tasks, i.e., presentation attack detection. Code and models are available at https://github.com/zhenglab/FFDBackbone.
@@ 1 Introduction
HUMAN faces, being unique and readily available biometric data, play a pivotal role in various research fields, including face recognition [108], generation [75], and analysis [88]. Unfortunately, facial information leakage and theft have spawned a clandestine industry where criminals exploit forged faces to perpetrate illegal activities, posing an extreme threat to national security, social stability, and individual well-being. Meanwhile, diverse GAN-based and Diffusion Model-based methods in Artificial Intelligence Generated Content (AIGC) have drastically reduced the barriers to creating lifelike forged faces. The proliferation of these forged faces online has led to significant crises in the security and authenticity of information. Therefore, distinguishing digitally forged faces created by image/video synthesis algorithms, i.e., Face Forgery Detection (FFD), also referred to as Deepfake detection, has become a critical issue in both research and industrial communities.
@@
Figure 1: Quantitative comparison of FFD methods using various backbones without pre-training (Unpre-trained), as well as pre-training with supervised learning (SL) or self-supervised learning (SSL) on ImageNet-1k (1k) or further on real face (Face) datasets. The average AUC across three cross-datasets (Celeb-DF, DFDC, and FFIW) better reflects the generalization of FFD models.
Unlike general image classification focusing on salient objects with specific structures and semantics, FFD requires identifying forgery cues that can appear irregularly anywhere on the face. As a result, many conventional CNN-based and Transformer-based architectures perform well on general tasks but often fail on FFD when trained end-to-end from scratch, as shown in the Unpre-trained results of Fig. 1, backbones without pre-training typically perform poorly and may fail to converge, whereas those pre-trained with supervised learning on ImageNet-1k (SL-1k) show significant improvement, similar to the findings in DeepfakeBench [125]. This marked difference is consistent across backbones including Xception [21], EfficientNet [99], and ViT [32], substantiating why most existing FFD methods [80, 140, 124, 30, 19, 121] employ pre-trained backbones as their foundational units. Meanwhile, SL-1k results also indicate that the performance of existing FFD methods such as CORE [80], CADDM [30], and UIA-ViT [140] primarily depends on their pre-trained backbones, which provide slight improvements over Xception, EfficientNet-b3, and ViT-B, respectively.
Due to the diverse forgery patterns generated by different synthesis algorithms (e.g., entire face synthesis [52], face swapping [71], face attribute editing [20], and face reenactment [84]), FFD models often overfit to specific forgery patterns in training datasets, limiting their generalization to unseen data [95, 124, 122]. Consequently, recent FFD research primarily leverages pre-trained backbones and focuses on designing advanced fine-tuning techniques, including data augmentation to expand forgery diversity [95, 62, 121], as well as attention mechanisms and feature decoupling to assist the backbone in better capturing spatial forgery cues [136, 33, 98, 140, 124]. However, these backbones are mostly adopted from previous FFD work or selected based on strong performance in general tasks [95, 140, 67], without a comprehensive exploration of which pre-training configurations improve FFD’s base performance. Although LipForensics [42] and RealForensics [41] utilize visual speech recognition and video-audio alignment as pre-training tasks, they depend on specific datasets with additional modality and are tailored to their particular approaches and applications.
Given the critical role of pre-trained backbones in FFD models, along with their extensive research in self-supervised learning (SSL) on general CV tasks [51, 116, 86], we argue that if SSL can be adapted to FFD and help to boost its performance. Thus, in this work, we systematically investigate pre-trained backbones with architectures such as ResNet and ViT via SSL on ImageNet-1k, including MoCo [45], SimSiam [16], MAE [44], and BEiT [83]. As shown by the SSL-1k results in Fig. 1, these backbones further improve FFD performance by benefiting from SSL’s enhancement of image representation. Notably, MoCo v3 [17] and BEiT v2 [83] boost ViT’s image feature representational capacity and achieve substantial gains over ViT with SL-1k. Therefore, we dive deep to integrate SSL pre-training into FFD workflow and tailor them to tackle FFD challenges, yielding SSL-pretrained backbones for enhancing FFD models’ sensitivity in identifying forgeries, as detailed in Sec. 3.2.
Indeed, facial images differ significantly from generic natural images due to their unique characteristics, including distinct shapes, keypoint distributions, and locally consistent displacement fields [47]. As synthesis techniques advance and fake faces become increasingly realistic, uncovering implicit forgery cues amid complex facial features becomes more challenging. However, existing FFD models primarily utilize backbones pre-trained on ImageNet [125, 121] or other natural image datasets such as LAION-2B [92] and WIT [85], which lack facial domain knowledge and may insufficiently represent the complex characteristics of facial images. It is worth noting that facial features encompass comprehensive information beyond semantic attributes such as identity or gender, as overfitting to these attributes can hinder generalization [124, 30]. Therefore, we further introduce real facial images to feed the SSL-pretrained backbones with face-specific knowledge, which can contribute to substantial improvements in FFD performance. As shown in the SSL-1k&Face results in Fig. 1, the backbones further pre-trained as FaceMAE, FaceMoCo, and FaceBEiT (FB) achieve varying degrees of improvement, with FB exhibiting particularly significant gains. Refer to Sec. 3.2 for more analysis.
In the above empirical studies of backbones with different pre-training configurations, we employ a basic end-to-end fine-tuning structure, i.e., Backbone and FC with Cross-Entropy loss, which is commonly used and directly reveals the backbone’s ability to support downstream task performance [86, 17, 44]. Moreover, in response to the subtle nature of forgery cues, we propose a competitive backbone fine-tuning framework that leverages dual backbones with an uncertainty-based fusion module to competitively identify implicit forgery cues, thereby improving the generalization performance of the pre-trained backbone. As shown by the SSL-1k&Face results in Fig. 1, our competitive fine-tuning method (CFB) achieves a $2.22\%$ significant performance improvement over its pre-trained backbone (FB), with additional experiments on other backbones presented in Table VII.
In practice, existing FFD methods typically adopt an empirical threshold of $0.5$ to convert continuous prediction probabilities ($p\in[0,1]$) into discrete labels (real or fake) [55]. However, this classification threshold depends on the probability distribution predicted by the FFD model on the test sets, and $0.5$ is not necessarily suitable. To address this, we propose a threshold optimization mechanism that integrates prediction probability and its confidence to compute a relatively optimal threshold, thereby enhancing the practicality of FFD models during inference.
@@
Additionally, to explore the potential and practicality of our FFD method, we validate it from digital face to physical forgery face detection, which is known as presentation attack detection (PAD)[34], face anti-spoofing[50, 114, 106, 73, 69, 59], and face liveness detection [9]. The PAD task focuses on identifying fake faces made from physical materials, such as printed photos, replayed images/videos, and 3D masks, all of which pose a significant threat in physical environments, particularly to face recognition systems.
@@
In summary, the main contributions of this work are as follows: (1) we systematically analyze the crucial role of pre-trained backbones in FFD task and develop an elaborate backbone that incorporates self-supervised learning and real faces to enhance facial feature representation for better forgery cue identification; (2) we build a competitive backbone fine-tuning framework with a decorrelation constraint and uncertainty-based fusion module to stimulate the backbone’s identification of generalization-benefiting implicit forgery cues; (3) we devise a threshold optimization mechanism using prediction probabilities and their uncertainty to calculate the optimal threshold from the seen test set, which improves discrimination accuracy on unseen test sets; (4) we conduct extensive experiments to demonstrate the superiority and potential of our FFD method, achieving significantly better generalization than previous methods in FFD and extra face-related PAD tasks.
@@ 2 Related Work
We start by summarizing existing FFD methods, then provide an overview of supervised and self-supervised learning for backbone pre-training.
@@ 2.1 Face Forgery Detection
Thanks to the creation of numerous face forgery datasets such as FaceForensic++ [91], Celeb-DF [63], and DFDC [29], deep learning-based FFD methods have made considerable advancements and can achieve outstanding accuracy within training datasets (exceeding 9595%). However, their performance drops sharply when encountering unseen forgery patterns across different datasets. Recent FFD work aims to improve model generalization by implementing specific fine-tuning methods such as data augmentation, attention mechanisms, forgery feature decoupling, and temporal inconsistency, as well as spatial and frequency fusion.
@@
Due to the presence of redundant forgery-independent features in forged faces (e.g., identity, natural artifacts), the distinction of forgery-dependent features is disrupted and hindered. Researchers leveraged disentangled representation learning [110] or game-theoretical approaches to enhance the backbone’s ability for extracting forgery clues. Specifically, disentangled representation learning-based FFD methods primarily adopted a training process that involved exchanging decoupled forgery-independent and forgery-dependent features to reconstruct pairs between real and forged faces, enhancing the discriminative performance of the classification branch with the backbone encoder [22, 48, 120, 5, 124]. Game-theoretical-based FFD methods [31, 128] utilized the knowledge of game theory to decouple unrelated regions from source and target images, promoting FFD performance with the support of Shapley value [94].
FTCN [137] and AltFreezing [112] explored temporal inconsistencies in forged videos. FTCN reduced the spatial convolution kernel, while AltFreezing alternated between spatial and temporal freezing, both aiming to encourage ResNet3D to explore generic features of forged face videos. Choi et al[19] employed StyleGRU and a style attention module to exploit the temporal low-variance properties of style latent vectors in each frame for detecting forged videos. TFCU [40] identifies spatially discriminative features in fine detail by unraveling temporal forgery cues. Additionally, Wang et al[107] and Miao et al[77] leveraged frequency domain information to assist the backbone in capturing more fine-grained artifact features.
Benefiting from the strong transfer capabilities of large-scale pre-trained models like CLIP in downstream tasks, recent studies have explored designing adapters tailored for FFD task. ForensicsAdapter [23] designs a CLIP adapter to transfer knowledge and enhance CLIP visual tokens for FFD generalization. Yan et al[126] proposed a lightweight spatio-temporal adapter for CLIP, effectively enabling spatio-temporal joint learning. M2F2-Det [38] employs face-forgery prompt learning with CLIP to enhance generalization and integrates an LLM to generate detailed textual explanations of its detection decisions. DFD-FCG [43] introduces a facial component guidance mechanism to improve the generalization of spatial learning by prompting the model to focus on key facial regions. RepDFD [64] reprograms CLIP model for FFD task by operating solely on its image and text inputs without tuning internal parameters. Additionally, UNITE [58] introduces an attention-diversity loss and employs SigLIP model for fully synthetic video detection. Zhang et al[134] introduce a text and image-aware feature alignment formulation to enhance multi-modal representation learning for Deepfake Detection VQA task.
After reviewing previous FFD methods and their implementations, we find that these methods consistently equip various fine-tuning techniques to backbones, such as ResNet [46], Xception [21], EfficientNet [99], and ViT [32], which are pre-trained on natural image datasets. These backbones are often used directly in the fine-tuning while overlooking the crucial roles of both the pre-training and inference, as shown in Fig. 3(top). In contrast, we endeavor to offer comprehensive research insights and guidance for FFD generalization by revisiting the complete FFD workflow.
@@ 2.2 Supervised and Self-Supervised Learning
Early backbone research utilized supervised learning (SL) with an image classification task as the pre-training approach, focusing on designing more sophisticated network architectures to improve backbone capacity and capability. During this period, numerous well-designed networks developed for the ImageNet Large-Scale Visual Recognition Challenge demonstrated outstanding performance, such as VGG, ResNet, and SENet. Benefiting from Transformers’ global contextual attention design, backbone network architectures have gradually transitioned from CNNs to Transformers. In particular, Vision Transformers (ViTs) [32] have come to dominate the field of computer vision (CV), achieving superior performance compared to ResNet, Xception, and EfficientNet.
@@
Given the significance of learning approaches with different pretext tasks for backbones in the pre-training stage, we conduct a comprehensive empirical study to investigate their performance in FFD task and provide valuable guidelines for their selection and optimization.
@@ 3.1 Research Overview
Our work leverages backbone pre-training as the starting point, and following our investigation of different pre-training configurations, we construct a comprehensive research overview, as shown in Fig. 2. This overview covers the complete FFD workflow: (1) the network architectures, (2) the learning approaches with SL and SSL, as well as the datasets such as ImageNet and real faces during pre-training, (3) the adaptation methods to downstream tasks during fine-tuning, and (4) the application mechanisms to maximize performance during inference.
@@
Figure 2: Overview of our research to revisit the complete FFD workflow from pre-training and fine-tuning to inference in discriminant results. This includes network architectures, datasets, learning approaches with supervised learning (SL) and self-supervised learning (SSL), which encompass contrastive learning (CL) and masked image modeling (MIM) pretext tasks, as well as our competitive backbone fine-tuning framework and threshold optimization mechanism.